# Plenoptic Cameras; Light Field; Calibration; Refocusing; Simulation

### 20 Feb 2024 v1

@@ 4.1 Light Field Decoding and SPC Calibration
First, the popular decoding and calibration theme presented by Dansereau et al. Dansereau et al. (2013) is revisited. In this work the raw plenoptic camera image is first de-multiplexed into an integer indexed two plane parametrization $L(i,j,k,l)$. These indices are then transformed into metric rays and propagated through the main lens. The combination of these steps yields an intrinsics matrix
@@
The relevant step with respect to the gap between the exit pupil and the principal plane in this process is the division of the integer indices by the respective spatial frequencies of the pixels and microlenses via the matrix $\mathbf{H}_{abs}^{\theta}$. As explained in section 2.1 the grid of MICs corresponds to the scaled grid of microlens centers. Accordingly, the sampling rate for the microlens plane has to be scaled down, or equivalently the pixel sampling rate has to be scaled up by the inverse factor. Dansereau et al. Dansereau et al. (2013) acknowledge this fact and choose the latter option, by introducing a scaling factor, which in our notation (compare section B.1) corresponds to $M_{\text{proj}}=\left(1+\frac{f_{m}}{d}\right)^{-1}$. This scaling, however, assumes a projection center at the main lens principal plane. Using the exit pupil instead, the correct rescaling is given by $M_{\text{proj}}=\left(1+\frac{f_{m}}{d-X}\right)^{-1}$ as visualized in figure 7.
@@
Similar cases of general parameters compensating for the model inaccuracies are presented by Monteiro et al. Monteiro et al. (2019) and Zhang et al. Zhang et al. (2018). Both make use of the same decoding process as Dansereau et al. (2013) and build upon the idea of directly relating the camera-side and scene light field. Monteiro et al. Monteiro et al. (2019) slightly reduce the intrinsics matrix shown above and subsequently use it in order to create an equivalent array of cameras for the scene-side light field. Zhang et al. Zhang et al. (2018) follow a similar approach by first relating the de-multiplexed light field in form of sub-aperture images with the scene-side, metric light field and basing further calculations on the latter. In both works, the interpretation of the intrinsic matrix parameters is not relevant, nor are these parameters used to directly reconstruct the main lens properties. Accordingly, the presented methods do not require a re-formulation using a more complex main lens model.
@@ 4.2 Depth Reconstruction
with system-dependent parameters $a_{0}=\frac{f_{m}\cdot d_{\text{ML}}}{s_{px}\cdot d}$, $a_{1}=\frac{\textbf{o}_{f}\cdot a_{0}}{f_{M}}$ and a shift parameter $\rho=\frac{s_{px}\cdot(F-F^{\prime})}{f_{m}\cdot d_{\text{ML}}}$. There are two problems in the deduction of this model. First, the shift parameter $\rho$ is not correctly deduced under the premise of light field data decoded by the method of Dansereau et al. Dansereau et al. (2013) and second, the exit pupil is ignored. In the following, a corrected version is presented, which also explains, how these two problems nearly neutralize each other and lead to the same general model, albeit with different parameter interpretations.
@@
The reason for this discrepancy is the implicitly used incorrect assumption in Pertuz et al. (2018), that the grid of microlens image centers equals the grid of microlens centers, i.e. $d_{\text{ML}}=d_{\text{MLI}}$. While a light field in general could be re-parametrized with this step size in the $ST$ plane, this requires exact knowledge of the camera and MLA geometry, which is usually unknown and not considered in the decoding process of Dansereau et al. (2013). Using the correct shift parameter for light field data following Dansereau et al. (2013) instead and rearranging the corresponding equation 10 results in
@@
Overall, the results of Pertuz et al. Pertuz et al. (2018) are not entirely incorrect, but the used light field representation based on Dansereau et al. (2013) is simply not matching the implicit assumptions used for the model deduction. More specifically, the main problem of Pertuz et al. (2018) is the definition of the shift parameter $\rho$ for light field data decoded similar to Dansereau et al. (2013), but based on known microlens centers instead of MICs. While the light field could theoretically be re-parametrized by using the parallel projections of the microlens centers onto the sensor, this would require exact knowledge of the SPC intrinsics, namely the MLA parameters as well as its placement relative to the main lens and sensor.
@@ 5.1 Simulation Environment
  • Simulation of aspheric lenses and zoom lenses
@@
This simulation facilitates a quick generation of a broad range of plenoptic cameras, such as the one exemplarily shown in figure 8, and is used in the following to validate the formal analysis of the sections 2.3, 3 and 4.2.
@@
Figure 8: Cross-section and rendering of an exemplary evaluation setup: A fully modeled lens with adjustable aperture and configurable lens group movements is combined with a two plane MLA model and a sensor in order simulate a plenoptic camera via ray tracing. Calibration patterns are placed at different distances in front of this setup to verify the analytical models. Note, that the housing of camera and lens were removed here for the purpose of visualization.
@@ 5.2 Experiments
For this validation the five lenses listed in table 1 were selected from the database Claff .
@@
While the first lens presents the ideal case of $X\approx 0mm$, i.e. the exit pupil coinciding with the camera-side principal plane, the remaining four lenses present interesting cases with varying relationships between $X$ and the lenses’ focal lengths $f_{M}$. Each of these lenses is used in two SPC configurations, one with a finite focus distance $\textbf{o}_{f}<\infty$ and another one focused at infinity, $\textbf{o}_{f}=\infty$. To this end, the MLA placement with respect to the main lens, i.e. the distance $d$ between the MLA and the camera-side principal plane of the main lens, can be calculated by the thin lens equation 1 as $d=\frac{f_{M}\cdot\textbf{o}_{f}}{\textbf{o}_{f}-f_{M}}$ for the finite case and simply be set to $d=f_{M}$ for $\textbf{o}_{f}=\infty$. The remaining microlens parameters are automatically initialized using the main lens and sensor properties to fulfill the following two constraints. First, the microlens f-number needs to match that of the main lens in order to optimally cover the sensor area Ng et al. (2005) and second, a predefined number of $129\times 129$ microlens images should be visible on the sensor in order to guarantee this resolution for the sub-aperture images. The resulting parameters are then fine-tuned by hand to accommodate the approximative nature of the f-number constraint and to guarantee that the MICs coincide with the centers of sensor pixels. This optimal MIC positioning has two effects: First, it renders the resampling during the decoding process of Dansereau et al. (2013) unnecessary. After normalizing the raw images with white images to get rid of vignetting effects Yu (2004), the sub-aperture images in our setups can directly be extracted by combining the same relative sensor pixels from each microlens image Ng et al. (2005). And second, the evaluation can concentrate on validating the refocusing itself instead of additionally dealing with the compensation of interpolation artifacts from the decoding process. The full setups can be found in appendix C and with each of these ten setups, five experiments are performed:
(I) MICs and Exit Pupil: The exit pupil as origin of the MICs is verified by first tracing ray bundles from the main lens aperture center through the main lens and MLA onto the sensor. This results in a set of sensor hits for every microlens. Due to the small variance within such a set, the mean is considered to represent the ground truh position of the microlens image center. In a second step, rays are traced from these sensor positions through the corresponding microlens centers and the convergence location of the resulting ray bundle is calculated in two ways. First by performing a line search along the optical axis for the minimum blur spot position of the ray bundle.
@@
(II) o and $S$: A calibration pattern, more specific a Siemens star with four spokes, is placed at various distances in front of the camera and after de-multiplexing the plenoptic camera image, the sub-aperture image shift $S$, which is required to focus onto the given target distance o, is measured. This is done via line search, i.e. by repeatedly refocusing the image with a simple shift-and-sum algorithm Ng et al. (2005) and calculating the sharpness of the refocused image. Here, the variance of the Laplacian Pertuz et al. (2013) is used as metric for image sharpness and the shift value with the highest score is considered the optimum. This procedure results in tuples of ground truth distances o and measured shifts $S$, which are then used to verify the connection $S(\textbf{o})$ as formally described in equation 6.
@@
Note, that instead of verifying $ERR_{\tilde{\textbf{o}}}$ and $ERR_{\tilde{S}}$, the equations 9 and 10 as well as the shift error model in 11 could also directly be validated with the data measured and calculated in the experiments (III) and (IV). However, these equations do not include a comparison between the incorrect estimations and the ground truth refocusing distances. Therefore, the indirect validation of those models by means of the resulting refocusing errors was preferred. (V) Validation of section 4.2: As analyzed in that section, the overall formulation of the model presented by Pertuz et al. Pertuz et al. (2018) is correct, but the model parameters have a different interpretation under the assumption of light field data decoded by the method of Dansereau et al. (2013). To verify the corrected model, the parameters $a_{0}$ and $a_{1}$ are first calculated based on the formula of Pertuz Pertuz et al. (2018), then according to our model from section 4.2, and finally fitted to the data of (II), i.e. the set of shift-distance-pairs with measured shifts and ground truth target distances. For this parameter fitting, a grid search for the best parameters by means of the RMSE was performed with a grid explicitly containing both directly calculated parameter sets.
@@ 5.3 Results and Discussion
The mean of the absolute differences $|S(\textbf{o})_{GT}-S(\textbf{o})_{measured}|$ between the measured and directly calculated ground truth shift values across all 10 setups is 0.008 px with a variance of 0.0026 px. The worst single setup is the finitely focused Olympus setup with a mean of absolute differences of 0.04 px and a variance of 0.025 px. These values are in the range of expected inaccuracies resulting from the image processing methods involved. Especially the interpolation steps required by the shift-and-sum refocusing Ng et al. (2005), but also the rather simple (de)focus measure, which is prone to interference errors, are limiting factors, which prevent a higher accuracy.
@@
A similar situation is observable in the inverse case, i.e. for the model $\textbf{o}(S)$ from equation 7, as presented in figure 12. Due to the wide range of target distances, in this case the relative absolute differences $\frac{|\textbf{o}(S)_{GT}-\textbf{o}(S)_{measured}|}{\textbf{o}(S)_{GT}}$ are calculated to quantify the results from figure 12. For the finite setups, the mean of these relative absolute differences is 0.13 % with a variance of $2\cdot 10^{-5}$, while in the infinite cases, the overall mean is 0.72 % with a variance of 0.01 %. Here, the infinitely focused Zeiss setup has the worst performance with a relative absolute difference mean of 0.9 % and a variance of 0.004 %. The worse performance of the infinitely focused setups is a consequence of the smaller range of shift values representing a larger refocusing range (compare figure 12) resulting in a greater susceptibility to small shift changes. Nevertheless, the overall performance confirms the model $\textbf{o}(S)$, again within the constraints posed by the involved image processing steps.
@@
Figure 13: Results of experiment (III): Relative object distance error resulting from a correctly estimated shift, but incorrect object distance estimation, i.e. the latter ignores the exit pupil. The thick lines indicate the predicted ground truth error based on equation 13 and equation 14 and the points indicate the measurements. Left: Error for the setups with finite focus. For a comparative visualization of the different target distance ranges, the results are shown based on the relative object distance $\lambda=\frac{\textbf{o}}{\textbf{o}_{f}}$. Right: Error for setups focused at infinity.
@@
(III) The results of the $ERR_{\tilde{\textbf{o}}}$ verification experiment are shown in figure 13. These graphs again show the error based on measurements compared to the directly calculated ground truth error based on the error models presented in section 3. The mean of absolute differences between the measured and calculated error values is 0.003 with a variance of $10^{-5}$, which validates our models limited only by the accuracy of image processing methods and the optical properties of the chosen lenses. More specific, the small deviations from the expected values, which are even present in the baseline case for the Rodenstock lens, can be explained as in experiment (II) by the interpolation operations required by the shift-and-sum refocusing algorithm Ng et al. (2005) for non-integer shift values. The larger deviation visible at close range for the Olympus lens with finite focus distance is a result of the optical limits of this lens which can be explained as follows. A single sub-aperture image contains one pixel per microlens. Hence, such an image can be considered sharp, if the calibration target is imaged onto the MLA with all blur spot sizes smaller than the microlens diameter $d_{\text{ML}}$. For the Olympus lens however, these blur spot sizes increase more drastically in the close range than those of the other lenses, leading to severe defocus blur even in the sub-aperture images. This in turn can produce interference artifacts during the refocusing and thereby lead to fluctuating contrast measurements. And since the refocus distance o is determined by these measurements, this affects the error calculated by $\frac{\textbf{o}-\tilde{\textbf{o}}(S(\textbf{o}))}{\textbf{o}}$ as in equation 13.

### 05 Apr 2024 v2

@@ 4 Revisiting SPC Methods
4.1 Equivalent Ray Model
First, the equivalence between the refocusing model in equation 7 and the ray intersection model presented by Hahne et al. Hahne et al. (2016) is proven. Instead of basing the model on decoding scheme of Dansereau et al. Dansereau et al. (2013) in Hahne et al. (2016) an approach building upon the intersection of chief rays is presented in order to calculate the refocusing distance for a resampling of the raw plenoptic camera image. A comprehensive notation transfer into our setup is given in section B.1.
Figure 7: Image formation (light blue) for an object point located at distance o from $H_{\text{scene}}$. The image of this point, located at distance i from $H_{\text{cam}}$, is seen by multiple microlenses and its projections onto the sensor have a metric disparity of $\hat{S}$. In order to determine the object distance o, Hahne et al. Hahne et al. (2016) propose to intersect ray functions (red) from two of the images and transfer the resulting image distance i to the scene via the thin lens equation.
The basic idea of Hahne et al. (2016) as depicted in figure 7 is the selection of two pixels on the sensor, which show scene points from the desired focus plane, and tracing rays from these through the respective microlens centers. The resulting camera-side intersection determines the distance of the virtual image inside the camera from the main lens and accordingly, the thin lens equation can be applied in order to calculate the corresponding object or refocusing distance. Without loss of generality, the following calculations will assume an MLA with one microlens center located on the main lens’ optical axis.
Given a sub-aperture image shift $S$ in pixel as in section 2.3, this translates to a metric pixel disparity $\hat{S}$ on the sensor by
$\hat{S}=-\frac{s_{px}}{S}$ (17)
with the sign flip resulting from the differing conventions used throughout this work and Hahne et al. (2016). Under the premise of a well configured plenoptic camera with a regular microlens grid, any two pixels from neighboring microlenses with a disparity of $\hat{S}$ can be chosen to calculate the image distance. To simplify the calculations, the first ray is chosen to run along the optical axis as shown in figure 7 and the second ray is based on the pixel at sensor position $d_{\text{MLI}}+\hat{S}$ going through the neighboring microlens center. According to Hahne et al. (2016) (compare section B.1) this leads to two ray functions
$f(z)=0\quad\text{and}\quad\tilde{f}(z)=\frac{d_{\text{ML}}-(d_{\text{MLI}}+\hat{S})}{f_{m}}\cdot z+d_{\text{ML}}$ (18)
which intersect at
$z_{i}=-\frac{d_{\text{ML}}\cdot f_{m}}{d_{\text{ML}}-(d_{\text{MLI}}+\hat{S})}.$ (19)
The crucial part, that sets Hahne et al. (2016) apart from the works reviewed in the following sections, is the correct calculation of the microlens image center distance $d_{\text{MLI}}$ based on the exit pupil (compare the calculation of $u_{c,j}$ in table 4) via
$d_{\text{MLI}}=\frac{d_{\text{ML}}}{d-X}\cdot f_{m}+d_{\text{ML}}$ (20)
which yields
$z_{i}$ $=-\frac{d_{\text{ML}}f_{m}}{d_{\text{ML}}-(\frac{d_{\text{ML}}}{d-X}\cdot f_{m}+d_{\text{ML}}+\hat{S})}$
$=\frac{(d-X)d_{\text{ML}}f_{m}}{(d-X)\hat{S}+f_{m}d_{\text{ML}}}.$ (21)
This intersection results in the image distance $\textbf{i}=d-z_{i}$ measured from $H_{\text{cam}}$ and can be used to calculate the object distance via the thin lens equation by
o $=\left(\frac{1}{f_{M}}-\frac{1}{\textbf{i}}\right)^{-1}=\left(\frac{1}{f_{M}}-\frac{1}{d-z_{i}}\right)^{-1}$
$=\frac{f_{M}\cdot d-f_{M}\cdot z_{i}}{(d-f_{M})-z_{i}}$
$=\frac{f_{M}\cdot d\cdot((d-X)\hat{S}+f_{m}d_{\text{ML}})-f_{M}\cdot(d-X)d_{\text{ML}}f_{m}}{(d-f_{M})\cdot((d-X)\hat{S}+f_{m}d_{\text{ML}})-(d-X)d_{\text{ML}}f_{m}}$
$=\frac{f_{M}\cdot d\cdot\frac{d-X}{f_{m}d_{\text{ML}}}\hat{S}+f_{M}X}{(d-f_{M})\cdot\frac{d-X}{f_{m}d_{\text{ML}}}\hat{S}-f_{M}+X}$
$=\frac{f_{M}\cdot(d\cdot\mathbf{\Delta}-S\cdot X)}{S\cdot(f_{M}-X)-\mathbf{\Delta}(f_{M}-d)}$ (22)
where equation 17 and the definition $\mathbf{\Delta}:=\frac{\Delta_{UV}}{\Delta_{ST}}=\frac{s_{px}(d-X)}{f_{m}d_{\text{ML}}}$ from section 2 are used in the last step. This equation equals the previously deduced equation 7 and thus proves the equivalence of both models.
@@ 4.2 Light Field Decoding and SPC Calibration
This section revisits the popular decoding and calibration theme presented by Dansereau et al. Dansereau et al. (2013). In that work the raw plenoptic camera image is first de-multiplexed into an integer indexed two plane parametrization $L(i,j,k,l)$. These indices are then transformed into metric rays and propagated through the main lens. The combination of these steps yields an intrinsics matrix
@@
The relevant step with respect to the gap between the exit pupil and the principal plane in this process is the division of the integer indices by the respective spatial frequencies of the pixels and microlenses via the matrix $\mathbf{H}_{abs}^{\theta}$. As explained in section 2.1 the grid of MICs corresponds to the scaled grid of microlens centers. Accordingly, the sampling rate for the microlens plane has to be scaled down, or equivalently the pixel sampling rate has to be scaled up by the inverse factor. Dansereau et al. Dansereau et al. (2013) acknowledge this fact and choose the second option, by introducing a scaling factor, which in our notation (compare section B.2) corresponds to
$M_{\text{proj}}=\left(1+\frac{f_{m}}{d}\right)^{-1}.$ (24)
This scaling, however, assumes a projection center at the main lens principal plane. Using the exit pupil instead, the correct rescaling is given by
$M_{\text{proj}}=\left(1+\frac{f_{m}}{d-X}\right)^{-1}$ (25)
as visualized in figure 8.
@@
Similar cases of general parameters compensating for the model inaccuracies are presented by Monteiro et al. Monteiro et al. (2019) and Zhang et al. Zhang et al. (2018). Both make use of the same decoding process as Dansereau et al. (2013) and build upon the idea of directly relating the camera-side and scene light field. Monteiro et al. Monteiro et al. (2019) slightly reduce the intrinsics matrix shown above and subsequently use it in order to create an equivalent array of cameras for the scene-side light field. Zhang et al. Zhang et al. (2018) follow a similar approach by first relating the de-multiplexed light field in form of sub-aperture images with the scene-side, metric light field, which they base all further calculations on. In both works, the interpretation of the intrinsic matrix parameters is not relevant, nor are these parameters used to directly reconstruct the main lens properties. Accordingly, the presented methods do not require a re-formulation using a more complex main lens model.
@@ 4.3 Depth Reconstruction
with system-dependent parameters
$a_{0}=\frac{f_{m}\cdot d_{\text{ML}}}{s_{px}\cdot d}\quad\text{and}\quad a_{1}=\frac{\textbf{o}_{f}\cdot a_{0}}{f_{M}}$ (27)
and a shift parameter
$\rho=\frac{s_{px}\cdot(F-F^{\prime})}{f_{m}\cdot d_{\text{ML}}}.$ (28)
There are two problems in the deduction of this model. First, the shift parameter $\rho$ is not correctly deduced under the premise of light field data decoded by the method of Dansereau et al. Dansereau et al. (2013) and second, the exit pupil is ignored. In the following, a corrected version is presented, which also explains, how these two problems nearly neutralize each other and lead to the same general model, albeit with different parameter interpretations.
@@
The reason for this discrepancy is the implicitly used incorrect assumption in Pertuz et al. (2018), that the grid of microlens image centers equals the grid of microlens centers, i.e. $d_{\text{ML}}=d_{\text{MLI}}$. While a light field in general could be reparametrized with this step size in the $ST$ plane, this requires exact knowledge of the camera and MLA geometry, which is usually unknown and not considered in the decoding process of Dansereau et al. (2013). Using the correct shift parameter for light field data following Dansereau et al. (2013) instead and rearranging the corresponding equation 10 results in
@@
Overall, the results of Pertuz et al. Pertuz et al. (2018) are not entirely incorrect, but the used light field representation based on Dansereau et al. (2013) is simply not matching the implicit assumptions used for the model deduction. More specifically, the main problem of Pertuz et al. (2018) is the definition of the shift parameter $\rho$ for light field data decoded similar to Dansereau et al. (2013), but based on known microlens centers instead of MICs. While the light field could theoretically be reparametrized by using the parallel projections of the microlens centers onto the sensor, this would require exact knowledge of the SPC intrinsics, namely the MLA parameters as well as its placement relative to the main lens and sensor.
@@ 5.1 Simulation Environment
  • Simulation of aspherical lenses and zoom lenses
@@
This simulation facilitates a quick generation of a broad range of plenoptic cameras, such as the one exemplarily shown in footnote 4, and is used in the following to validate the formal analysis of the sections 2.3, 3 and 4.3.
@@
Figure 9: Cross-section and rendering of an exemplary evaluation setup: A fully modeled lens is combined with a two plane MLA model and a sensor in order simulate a plenoptic camera via ray tracing. Calibration patterns are placed at different distances in front of this setup to verify the analytical models. Note, that the housing of camera and lens were removed for the purpose of visualization and the ISO 12233 pattern [^4] https://www.graphics.cornell.edu/~westin/misc/res-chart.html is used with permission of Cornell University.
@@ 5.2 Experiments
For the validation the five lenses listed in table 1 were selected from the database Claff .
@@
While the first lens presents the ideal case of $X\approx 0mm$, i.e. the exit pupil coinciding with the camera-side principal plane, the remaining four lenses present interesting cases with varying relationships between $X$ and the lenses’ focal lengths $f_{M}$. Each of these lenses is used in two SPC configurations, one with a finite focus distance $\textbf{o}_{f}<\infty$ and another one focused at infinity, $\textbf{o}_{f}=\infty$. To this end, the MLA placement with respect to the main lens, i.e. the distance $d$ between the MLA and the camera-side principal plane of the main lens, can be calculated by the thin lens equation 1 as $d=\frac{f_{M}\cdot\textbf{o}_{f}}{\textbf{o}_{f}-f_{M}}$ for the finite case and simply be set to $d=f_{M}$ for $\textbf{o}_{f}=\infty$. The remaining microlens parameters are automatically initialized using the main lens and sensor properties to fulfill the following two constraints. First, the microlens f-number needs to match that of the main lens in order to optimally cover the sensor area Ng et al. (2005) and second, a predefined number of $129\times 129$ microlens images should be visible on the sensor in order to guarantee this resolution for the sub-aperture images. The resulting parameters are then fine-tuned by hand to accommodate the approximating nature of the f-number constraint and to guarantee that the MICs coincide with the centers of sensor pixels. This optimal MIC positioning has two effects: First, it renders the resampling during the decoding process of Dansereau et al. (2013) unnecessary. After normalizing the raw images with white images to get rid of vignetting effects Yu (2004), the sub-aperture images in our setups can directly be extracted by combining the same relative sensor pixels from each microlens image Ng et al. (2005). And second, the evaluation can concentrate on validating the refocusing itself instead of additionally dealing with the compensation of interpolation artifacts from the decoding process. The full setups can be found in appendix D and with each of these ten setups, five experiments are performed:
(I) MICs and Exit Pupil: The exit pupil as origin of the MICs is verified by first tracing ray bundles from the main lens aperture center through the main lens and MLA onto the sensor. This results in a set of sensor hits for every microlens. Due to the small variance within such a set, the mean is considered to represent the ground truth position of the microlens image center. In a second step, rays are traced from these sensor positions through the corresponding microlens centers and the convergence location of the resulting ray bundle is calculated in two ways. First by performing a line search along the optical axis for the minimum blur spot position of the ray bundle.
@@
(II) o and $S$: A calibration pattern, more specific a Siemens star with four spokes, is placed at various distances in front of the camera and after demultiplexing the plenoptic camera image, the sub-aperture image shift $S$, which is required to focus onto the given target distance o, is measured. This is done via line search, i.e. by repeatedly refocusing the image with a simple shift-and-sum algorithm Ng et al. (2005) and calculating the sharpness of the refocused image. Here, the variance of the Laplacian Pertuz et al. (2013) is used as metric for image sharpness and the shift value with the highest score is considered the optimum. This procedure results in tuples of ground truth distances o and measured shifts $S$, which are then used to verify the connection $S(\textbf{o})$ as formally described in equation 6.
@@
Note, that instead of verifying $ERR_{\tilde{\textbf{o}}}$ and $ERR_{\tilde{S}}$, the equations 9 and 10 as well as the shift error model in equation 11 could also directly be validated with the data measured and calculated in the experiments (III) and (IV). However, these equations do not include a comparison between the incorrect estimations and the ground truth refocusing distances. Therefore, the indirect validation of those models by means of the resulting refocusing errors was preferred. (V) Validation of section 4.3: As analyzed in that section, the overall formulation of the model presented by Pertuz et al. Pertuz et al. (2018) is correct, but the model parameters have a different interpretation under the assumption of light field data decoded by the method of Dansereau et al. (2013). To verify the corrected model, the parameters $a_{0}$ and $a_{1}$ are first calculated based on the formula of Pertuz Pertuz et al. (2018), then according to our model from section 4.3, and finally fitted to the data of (II), i.e. the set of shift-distance-pairs with measured shifts and ground truth target distances. For this parameter fitting, a grid search for the best parameters by means of the RMSE was performed with a grid explicitly containing both directly calculated parameter sets.
@@ 5.3 Results and Discussion
The mean of the absolute differences $|S(\textbf{o})_{measured}-S(\textbf{o})_{GT}|$ between the measured and directly calculated ground truth shift values across all 10 setups is 0.008 px with a variance of 0.0026 px. The worst single setup is the finitely focused Olympus setup with a mean of absolute differences of 0.04 px and a variance of 0.025 px. These values are in the range of expected inaccuracies resulting from the image processing methods involved. Especially the interpolation steps required by the shift-and-sum refocusing Ng et al. (2005), but also the rather simple (de)focus measure, which is prone to interference errors, are limiting factors, which prevent a higher accuracy.
@@
A similar situation is observable in the inverse case, i.e. for the model $\textbf{o}(S)$ from equation 7, as presented in figure 13. Due to the wide range of target distances, in this case the relative absolute differences $\frac{|\textbf{o}(S)_{measured}-\textbf{o}(S)_{GT}|}{\textbf{o}(S)_{GT}}$ are calculated to quantify the results from figure 13. For the finite setups, the mean of these relative absolute differences is 0.13 % with a variance of $2\cdot 10^{-5}$, while in the infinite cases, the overall mean is 0.72 % with a variance of 0.01 %. Here, the infinitely focused Zeiss setup has the worst performance with a relative absolute difference mean of 0.9 % and a variance of 0.004 %. The worse performance of the infinitely focused setups is a consequence of the smaller range of shift values representing a larger refocusing range (compare figure 13) resulting in a greater susceptibility to small shift changes. Nevertheless, the overall performance confirms the model $\textbf{o}(S)$, again within the constraints posed by the involved image processing steps.
@@
Figure 14: Results of experiment (III): Relative object distance error resulting from a correctly estimated shift, but incorrect object distance estimation based on the assumption $X=0$. The thick lines indicate the predicted ground truth error based on equation 13 and equation 14 and the points indicate the measurements. Left: Error for the setups with finite focus. For a comparative visualization of the different target distance ranges, the results are shown based on the relative object distance $\lambda=\frac{\textbf{o}}{\textbf{o}_{f}}$. Right: Error for setups focused at infinity.
@@
(III) The results of the $ERR_{\tilde{\textbf{o}}}$ verification experiment are shown in figure 14. These graphs again show the error based on measurements compared to the directly calculated ground truth error based on the error models presented in section 3. The mean of absolute differences between the measured and calculated error values is 0.003 with a variance of $10^{-5}$, which validates our models limited only by the accuracy of image processing methods and the optical properties of the chosen lenses. More specific, the small deviations from the expected values, which are even present in the baseline case for the Rodenstock lens, can be explained as in experiment (II) by the interpolation operations required by the shift-and-sum refocusing algorithm Ng et al. (2005) for non-integer shift values. The larger deviation visible at close range for the Olympus lens with finite focus distance is a result of the optical limits of this lens which can be explained as follows. A single sub-aperture image contains one pixel per microlens. Hence, such an image can be considered sharp, if the calibration target is imaged onto the MLA with all blur spot sizes smaller than the microlens diameter $d_{\text{ML}}$. For the Olympus lens however, these blur spot sizes increase more drastically in the close range than those of the other lenses, leading to severe defocus blur even in the sub-aperture images. This in turn can produce interference artifacts during the refocusing and thereby lead to fluctuating contrast measurements. And since the refocus distance o is determined by these measurements, this affects the error calculated by $\frac{\tilde{\textbf{o}}(S(\textbf{o}))-\textbf{o}}{\textbf{o}}$ as in equation 13.